For the random variable X defined by the probability function below, answer the following questions.

x 2 3 4 5 6 7

f(x) 0.25 0.15 0.10 0.15 0.10 0.25

a. Find the probability that X is even or at least 7.

b. Find and sketch the cdf.

c. Find the expected value and variance of X.

d. Find the expected value and variance of Y=2X+3.


1
Expert's answer
2021-01-26T04:44:11-0500

a.

"P(even\ \cup\ \geq7)=0.25+0.10+0.1+0.25=0.7"

b.


"F(2)=f(2)=0.25"

"F(3)=f(2)+f(3)=0.25+0.15=0.4"

"F(4)=f(2)+f(3)+f(4)=0.25+0.15+0.10"

"=0.5"

"F(5)=f(2)+f(3)+f(4)+f(5)"

"=0.25+0.15+0.10+0.15=0.65"

"F(6)=f(2)+f(3)+f(4)+f(5)+f(6)"

"=0.25+0.15+0.10+0.15+0.10=0.75"


"F(7)=f(2)+f(3)+f(4)+f(5)+f(6)+f(7)"

"=0.25+0.15+0.10+0.15+0.10+0.25=1"

Hence


"F(x) = \begin{cases}
 0, &\text{for } x<2 \\
 0.25, &\text{for } 2\leq x<3 \\
0.4, &\text{for } 3\leq x<4 \\
0.5, &\text{for } 4\leq x<5 \\
0.65, &\text{for } 5\leq x<6 \\
0.75, &\text{for } 6\leq x<7 \\
1, &\text{for } x\geq7 \\
\end{cases}"


c.


"E(X)=2(0.25)+3(0.15)+4(0.10)+5(0.15)"

"+6(0.10)+7(0.25)=4.45"

"E(X^2)=2^2(0.25)+3^2(0.15)+4^2(0.10)"

"+5^2(0.15)+6^2(0.10)+7^2(0.25)=23.55"

"Var(X)=E(X^2)-(E(X))^2=23.55-(4.45)^2"

"=3.7475"

d. "Y=2X+3"


"E(Y)=E(2X+3)=2E(X)+3"

"=2(4.45)+3=11.9"

"Var(Y)=Var(2X+3)=2^2Var(X)"

"=4(3.7475)=14.99"